# Radio-frequency assisted switching in perpendicular magnetic tunnel junctions

### v1

@@ Abstract
Spin-transfer torque magnetic random-access memory (STT-MRAM) relies on nanoscale magnetic tunnel junctions (MTJs) as its fundamental building blocks. Next-generation STT-MRAM requires strategies that simultaneously improve switching energy efficiency and device endurance. Here, we present the first study of perpendicular STT-MRAM writing assisted by radio-frequency (RF) spin torque. We show that applying a small-amplitude RF pulse prior to a direct-current (DC) writing pulse enhances the MTJ switching probability, with the efficiency gain increasing at lower RF frequencies. This RF+DC writing scheme enables shorter DC pulses, thereby improving device endurance. Analytical and numerical modeling qualitatively reproduces the experimental trends, while quantitative discrepancies indicate that realistic MTJ properties beyond idealized models play an important role in RF-assisted switching.
@@ Introduction
STT-MRAM is a leading candidate for next-generation nonvolatile memory, combining sub-nanosecond switching [1, 2, 3, 4, 5], CMOS compatibility [6, 7, 8, 9], scalability to future technology nodes [10], and radiation hardness [11]. Significant progress has been made toward commercialization, yet the technology is currently limited to niche applications [12]. Advancing the technology to mainstream implementation requires overcoming the key challenges of relatively low energy efficiency, high write current density, and limited endurance [8]. In perpendicular MTJs (p-MTJs), which constitute the core of contemporary STT-MRAM, the write current density must be kept high enough to guarantee reliable switching, but this comes at the cost of increased power consumption and the risk of tunnel barrier breakdown [13]. Scaling the MTJ size below 20 nm further exacerbates this trade-off by making the thermal stability and write reliability harder to balance [14].
The relatively high voltage pulses applied during the write operation accelerate degradation of the ultrathin MgO barrier, leading to reduced device lifetimes. This reliability challenge stresses the need for writing schemes that deliver switching at reduced voltage levels, without compromising retention or write error rates. An alternative, spin-orbit torque MRAM (SOT-MRAM), addresses the problem of dielectric breakdown by separating the read and write current paths [15, 16, 17, 18, 19, 20, 21]. Although this decoupling enables fast switching without stressing the MgO barrier, the three-terminal geometry of SOT-MRAM imposes a large footprint that constrains storage density and integration. Therefore, innovation in the engineering of the STT-MRAM stack, device design, and write protocols is crucial for the development of energy-efficient STT-MRAM.
A straightforward strategy to reduce the write voltage in STT-MRAM is to increase the duration of the DC write pulse. However, this approach relying on thermally-assisted switching comes at the cost of higher write energy and reduced energy efficiency [22, 23, 24]. The tradeoff between the write voltage amplitude and energy efficiency motivates the search for alternative write schemes [25] that achieve reliable switching without compromising endurance or efficiency.
An attractive approach is to combine the DC write pulse with an RF current pulse [26, 27, 28] – a method inspired by studies of microwave-assisted magnetization reversal in magnetic media [29, 30, 31, 32, 33, 34]. This RF-assisted strategy has previously been studied in in-plane MTJs and in all-metallic spin valves, where an enhanced switching probability was observed when the RF drive frequency was tuned to the ferromagnetic resonance (FMR) of the free layer (FL) [35, 36]. In thermally unstable, superparamagnetic in-plane MTJs, a strong enhancement of random telegraph switching has been demonstrated at sub-GHz frequencies, where RF excitation drives low-dimensional chaotic dynamics of the FL [37].
@@
Figure 2a illustrates the pulse sequence used for measurements of p-MTJ switching probability $\mathcal{P}$ ($0\leq\mathcal{P}\leq 1$) in the conventional DC write pulse switching regime. In this paper, we study switching from the low-resistance parallel (P) to the high-resistance antiparallel (AP) state at room temperature in zero external magnetic field. A large-amplitude square DC write pulse is applied to the device in the P state, followed by the resultant resistance readout and a longer reset pulse returning the MTJ back to the initial P state. The measurement is repeated $10^{5}$ times to reliably measure $\mathcal{P}$ for a given write pulse duration $t_{\mathrm{DC}}$ and amplitude $V_{\mathrm{DC}}$. Figure 2b shows the results of our measurements for a $D=45$ nm device as a function of $t_{\mathrm{DC}}$ for three values of $V_{\mathrm{DC}}$. All voltage values reported in this paper are voltages applied across the MTJ (Supplementary Note 1).
For studies of RF-assisted switching, the DC write pulse is preceded by an RF pulse of duration $t_{\mathrm{RF}}$, frequency $f_{\mathrm{RF}}$ and root mean square (RMS) amplitude $V_{\mathrm{RF}}$ as shown in Fig. 2c. A variable delay $\tau$ is introduced between the end of the RF pulse and the onset of the DC pulse; negative delay $\tau<0$ corresponds to temporal overlap of the RF and DC pulses. We study both the $\tau\geq 0$ and $\tau<0$ regimes of RF-assisted switching. The phase of the RF pulse in our measurements is random (Supplementary Note 1 and Supplementary Figure S1).
@@
Figure 5: RF-assisted increase in switching probability $\Delta\mathcal{P}$ as a function of RF frequency.
Figure 5 shows the dependence of $\Delta\mathcal{P}$ on $f_{\mathrm{RF}}$ for the $D=45$ nm device at $P_{\mathrm{RF}}=-21$ dBm and $\tau=-2$ ns. These data, taken at frequencies well below the FL FMR frequency (12 GHz, Supplementary Note 2), reveal that RF-assisted switching becomes more efficient with decreasing RF frequency. A qualitatively similar dependence of $\Delta\mathcal{P}$ on $f_{\mathrm{RF}}$ is observed in the $\tau\geq 0$ regime.
Theory
In order to understand the mechanisms of the experimentally observed RF-assisted switching, we made room-temperature ($T=300$ K) numerical macrospin and micromagnetic simulations of the switching process. Simulations were made for all the MTJ diameters considered in the experiments (Supplementary Note 3). Here we show the results for the device with $D=45$ nm. Simulations of an ensemble of multiple replicas were performed to compute the switching probability. The magnetization dynamics of each replica were computed using both micromagnetic and macrospin solvers (see Supplementary Note 3). For the micromagnetic analysis, $10^{2}$ replicas were simulated, while for the macrospin simulations $10^{4}$ replicas were used. The switching probability was then determined by counting the number of replicas that showed P to AP switching.
For these simulations, we first select the values of $V_{\mathrm{DC}}$ and $t_{\mathrm{DC}}$ that produce $\mathcal{P}=0.5$ in the absence of an RF pulse for each device. We then add an RF pulse preceding the DC pulse and calculat the increase in the switching probability $\Delta\mathcal{P}$ above the $\mathcal{P}=0.5$ baseline.
@@
Figure 6 shows the simulated $\Delta\mathcal{P}$ due to the RF pulse assist versus normalized RF power $\tilde{P}_{\mathrm{RF}}$ and delay $\tau$ between the RF and DC write pulses. The RF power is normalized to the minimal theoretical power required to switch the magnetic state with a DC current in the absence of thermal fluctuations ($T=0$ K) [42]. This minimal power is analytically calculated in the macrospin approximation for the values of the FL dimensions, the uniaxial magnetocrystalline anisotropy constant $K_{\mathrm{AN}}$, saturation magnetization $M_{\mathrm{s}}$, and Gilbert damping constant $\alpha$ given in Table I of Methods. Figure 6a shows macrospin $\Delta\mathcal{P}$ as a function of $\tilde{P}_{\mathrm{RF}}$ and $\tau$ when the phase constant $\phi$ of the RF voltage $\sqrt{2}V_{\mathrm{RF}}\cos(\omega_{\mathrm{RF}}t+\phi)$ is set to zero. Notice that $t=0$ corresponds to the time instant where the DC pulse is applied and $\omega_{\mathrm{RF}}=2\pi f_{\mathrm{RF}}$. These results are significantly different from the experiment in Fig. 3c and Fig. 4c. This is most clearly seen from the comparison of the inset in Fig. 6a showing $\Delta\mathcal{P}$ versus $\tau$ at $\tilde{P}_{\mathrm{RF}}=-5$ dB and the experimental data in Fig. 4c. Indeed, the measured $\Delta\mathcal{P}$ is a monotonically decreasing function of $\tau$, while the simulated $\Delta\mathcal{P}(\tau)$ has a peak at $\tau=-3$ ns. The micromagnetic results in Fig. 6b are similar to the macrospin with a slightly enhanced $\Delta\mathcal{P}$. The number of replicas in the macrospin simulations is large enough to yield a $\sim$1 % error in $\Delta\mathcal{P}$ while the micromagnetic results are precise within $\sim$10 %.
@@
The discrepancy with the experiment is expected because the simulations assume a fixed phase $\phi=0$, while the measurements average over random phases. To address this discrepancy, we performed macrospin simulations with $\phi$ uniformly distributed over $\pi/2,\pi/2]$ (Fig. 6c) and $\pi,\pi]$ (Fig. 6d) intervals, the latter directly mimicking the experimental conditions. The inset in Fig. 6d is qualitatively similar to the corresponding experimental results in Fig. 4c, although quantitative discrepancies remain. Similar results were obtained for devices with other studied MTJ diameters (Supplementary Note 3 and Supplementary Figure S3).
@@
(ii) prior to the RF+DC waveform, the distribution of replicas is assumed to be described by the Néel-Brown model [45] in the high-energy-barrier case. Therefore, the initial probability density function (PDF) of the replicas $p_{0}(m_{z})$ is given by the following expression:
$p_{0}(m_{z})=\frac{1}{Z}\exp\left(\frac{k_{\mathrm{eff}}}{2\tilde{T}}m_{z}^{2}\right)\,,\,m_{z}\in[0,1]\,$, (2)
where
[…]
is the partition function, $m_{z}$ is the out-of-plane component of the FL magnetization unit vector $\bm{m}=\bm{M}/M_{s}(T)$, with $M_{s}(T)$ the saturation magnetization at a certain temperature $T$, and $\mathrm{erfi}(.)$ is the imaginary error function. $k_{\mathrm{eff}}$ is the normalized effective magnetic anisotropy constant of the FL given by $k_{\mathrm{eff}}=2K_{\mathrm{AN}}/(\mu_{0}M_{s}^{2})-(D_{\mathrm{z}}-D_{\mathrm{\perp}})$. Here, $K_{\mathrm{AN}}$ is the uniaxial magnetocrystalline anisotropy constant in SI units, $D_{\mathrm{z}}$ and $D_{\mathrm{\perp}}$ are the demagnetizing factors along the symmetry axis and the perpendicular direction $(2D_{\perp}+D_{z}=1)$, and $\mu_{0}$ is the vacuum magnetic permeability. $\tilde{T}=k_{\mathrm{B}}T/(\mu_{0}M_{\mathrm{s}}^{2}V)$ is the normalized temperature, where $k_{\mathrm{B}}$ is the Boltzmann constant and $V$ is the FL volume.
@@
$\overline{V}_{\mathrm{ST}}=V_{\mathrm{DC}}+\frac{\sqrt{2}V_{\mathrm{RF}}}{\omega_{\mathrm{RF}}t_{\mathrm{DC}}}$ $\Big[\sin(\omega_{\mathrm{RF}}(t_{\mathrm{RF}}-\tau)+\phi)-\sin(\omega_{\mathrm{RF}}t_{\mathrm{RF}}+\phi)\Big]\,$, (4)
which is $V_{\mathrm{DC}}$ plus the average RF voltage over the duration of the DC pulse, with $t_{\mathrm{RF}}$ the duration of the RF pulse. The theoretical foundations of these five assumptions are discussed in Supplementary Note 3.
The PDF of the FL magnetization at a time $t$ is:
$p(m_{z},t)=\int_{0}^{1}p_{0}(m_{z0})\,p(m_{z},t|m_{z0},0)\,dm_{z0}\,$, (5)
where $m_{z0}$ is $m_{z}$ at $t=0$, which marks the onset of the $\overline{V}_{\mathrm{ST}}$ write pulse.
The magnetization switching dynamics is a damping-switching process in a uniaxial, bistable magnet between the two stable states $m_{z}=\pm 1$. In this case the magnetization trajectory from $m_{z0}$ to $m_{z}$ under the action of $\overline{H}=\overline{V}_{\mathrm{ST}}/V_{\mathrm{c0}}=\overline{\beta}_{\mathrm{ST}}/(\alpha k_{\mathrm{eff}})$ can be expressed in closed form by the following [44]: $q(m_{z},m_{z0},\overline{H})=\alpha k_{\mathrm{eff}}t$. The quantity $\overline{H}$ is the effective DC voltage pulse amplitude normalized to $V_{\mathrm{c0}}$, which is the critical DC voltage for switching at $T=0$. For the numerical integration of the magnetization dynamics, the dimensionless LLG equation is considered [43]. Within it, the dimensionless voltage is $\beta_{\mathrm{ST}}$, while the dimensionless critical voltage at zero temperature is given by $\alpha k_{\mathrm{eff}}$.
Solving the relation defining $q$ for $m_{z}$, we obtain: $m_{z}(t)=\mu_{z}(m_{z0},t,\overline{H})$, which allows us to rewrite Eq. 5 in the following form:
$p(m_{z},t)=\frac{p_{0}(\mu_{z}^{-1}(m_{z}))}{\lambda_{z}(m_{z})}\,$, (6)
where $\lambda_{z}=d\mu_{z}/dm_{z0}$.
We assume that the FL switches if $m_{z}$ changes sign compared to $m_{z0}$ after the write pulse. Starting with the initial distribution with $m_{z0}>0$, the switching probability can be computed by integrating the PDF over the range $m_{z}\in[-1,0]$ that defines the well of the switched magnetic state $m_{z}=-1$. Such a state cannot be reached by a deterministic dynamics in a time window of finite duration. The closest state to $m_{z}=-1$ reachable in a finite time window is $m_{z}=\mu_{z}(0,t,\overline{H})$, which is the final state of a deterministic trajectory in the time interval $[0,t]$ starting from the initial state $m_{z}(0)=0$. With these considerations, the switching probability $\mathcal{P}$ at $t$ is given by:
@@
$\mathcal{P}(t)=$ $\int_{\mu_{z}(0,t,\overline{H})}^{0}p(m_{z},t)\,dm_{z}=\int_{0}^{\mu_{z}^{-1}(0,t,\overline{H})}p_{0}(m_{z0})\,dm_{z0}\,=\frac{\mathrm{erfi}\left(\sqrt{\frac{k_{\mathrm{eff}}}{2\tilde{T}}}\mu_{z}^{-1}(0,t,\overline{H})\right)}{\mathrm{erfi}\left(\sqrt{\frac{k_{\mathrm{eff}}}{2\tilde{T}}}\right)}\,$. (7)
Taking into account that PDF depends on $\phi$ via $\overline{H}$, the switching probability for $\phi$ uniformly distributed over the experimentally relevant interval $\pi,\pi]$ is:
$P(t)=\frac{1}{2\pi\,\mathrm{erfi}\left(\sqrt{\frac{k_{\mathrm{eff}}}{2\tilde{T}}}\right)}\int_{0}^{2\pi}\mathrm{erfi}\left(\sqrt{\frac{k_{\mathrm{eff}}}{2\tilde{T}}}\mu_{z}^{-1}(0,t,\overline{H})\right)d\phi\,.$ (8)
@@
In the case of the fully random phase $\phi$ distributed in $\phi\in[-\pi,\pi]$ (Fig. 7c), analytical calculations do not predict an increase in $\mathcal{P}$, while numerical simulations show $\Delta\mathcal{P}>0$ for large values of $P_{\mathrm{RF}}$ and $\tau<0$. This is because $\mathcal{P}(\phi)$ in Eq. 7 is a periodic function of $\phi$, and the enhancement of $\mathcal{P}$ achieved for a given $\phi^{*}$ is approximately canceled by the reduction of $\mathcal{P}$ for the phase $\phi^{*}+\pi$. Although this cancellation is not exact, $\Delta\mathcal{P}$ resulting from the averaging in $\phi$ is small $\sim 0.01$.
In numerical simulations with sufficiently high $\tilde{P}_{\mathrm{RF}}$ and $\tau<0$, $\Delta\mathcal{P}$ is not negligibly small because a larger number of switching events with respect to those predicted by the analytical model occur in the time interval where the RF pulse adds constructively to the DC pulse voltage. This fact is not surprising since the assumption (v) of the analytical model does not hold anymore (Supplemental Note 3). As a result, numerical simulations produce higher $\Delta\mathcal{P}$ than the analytic model.

### v2

@@ Abstract
Abstract: Spin-transfer torque magnetic random-access memory (STT-MRAM) relies on nanoscale magnetic tunnel junctions (MTJs) as its fundamental building blocks. Next-generation STT-MRAM requires strategies that simultaneously improve switching energy efficiency and device endurance. Here, we present the first study of perpendicular STT-MRAM writing assisted by radio-frequency (RF) spin torque. We show that applying a small-amplitude RF pulse prior to a direct-current (DC) writing pulse enhances the MTJ switching probability, with the efficiency gain increasing at lower RF frequencies. This RF+DC writing scheme enables shorter DC pulses, thereby improving device endurance. Analytical and numerical modeling qualitatively reproduces the experimental trends, while quantitative discrepancies indicate that realistic MTJ properties beyond idealized models play an important role in RF-assisted switching.
@@ Introduction
STT-MRAM is a leading candidate for next-generation nonvolatile memory, combining sub-nanosecond switching [37, 26, 13, 54], CMOS compatibility [35, 2, 53, 22], scalability to future technology nodes [51], and radiation hardness [31]. Significant progress has been made toward commercialization, yet the technology is currently limited to niche applications [40]. Advancing the technology to mainstream implementation requires overcoming the key challenges of relatively low energy efficiency, high write current density, and limited endurance [53]. In perpendicular MTJs (p-MTJs), which constitute the core of contemporary STT-MRAM, the write current density must be kept high enough to guarantee reliable switching, but this comes at the cost of increased power consumption and the risk of tunnel barrier breakdown [56]. Scaling the MTJ size below 20 nm further exacerbates this trade-off by making the thermal stability and write reliability harder to balance [33].
The relatively high voltage pulses applied during the write operation accelerate degradation of the ultrathin MgO barrier, leading to reduced device lifetimes. This reliability challenge stresses the need for writing schemes that deliver switching at reduced voltage levels, without compromising retention or write error rates. An alternative, spin-orbit torque MRAM (SOT-MRAM), addresses the problem of dielectric breakdown by separating the read and write current paths [30, 24, 55, 25]. Although this decoupling enables fast switching without stressing the MgO barrier, the three-terminal geometry of SOT-MRAM imposes a large footprint that constrains storage density and integration. Therefore, innovation in the engineering of the STT-MRAM stack, device design, and write protocols is crucial for the development of energy-efficient STT-MRAM.
A straightforward strategy to reduce the write voltage in STT-MRAM is to increase the duration of the DC write pulse. However, this approach relying on thermally-assisted switching comes at the cost of higher write energy and reduced energy efficiency [3, 46]. The tradeoff between the write voltage amplitude and energy efficiency motivates the search for alternative write schemes [23] that achieve reliable switching without compromising endurance or efficiency.
An attractive approach is to combine the DC write pulse with an RF current pulse [17, 11] – a method inspired by studies of microwave-assisted magnetization reversal in magnetic media [49, 52]. This RF-assisted strategy has previously been studied in in-plane MTJs and in all-metallic spin valves, where an enhanced switching probability was observed when the RF drive frequency was tuned to the ferromagnetic resonance (FMR) of the free layer (FL) [14, 18]. In thermally unstable, superparamagnetic in-plane MTJs, a strong enhancement of random telegraph switching has been demonstrated at sub-GHz frequencies, where RF excitation drives low-dimensional chaotic dynamics of the FL [32].
@@ Results
RF-assisted switching measurements Figure 2a illustrates the pulse sequence used for measurements of p-MTJ switching probability $\mathcal{P}$ ($0\leq\mathcal{P}\leq 1$) in the conventional DC write pulse switching regime. In this paper, we study switching from the low-resistance parallel (P) to the high-resistance antiparallel (AP) state at room temperature in zero external magnetic field. A large-amplitude square DC write pulse is applied to the device in the P state, followed by the resultant resistance readout and a longer reset pulse returning the MTJ back to the initial P state. The measurement is repeated $10^{5}$ times to reliably measure $\mathcal{P}$ for a given write pulse duration $t_{\mathrm{DC}}$ and amplitude $V_{\mathrm{DC}}$. Figure 2b shows the results of our measurements for a $D=45$ nm device as a function of $t_{\mathrm{DC}}$ for three values of $V_{\mathrm{DC}}$. All voltage values reported in this paper are voltages applied across the MTJ (Supplementary Note 1).
For studies of RF-assisted switching, the DC write pulse is preceded by an RF pulse of duration $t_{\mathrm{RF}}$, frequency $f_{\mathrm{RF}}$ and root mean square (RMS) amplitude $V_{\mathrm{RF}}$ as shown in Fig. 2c. A variable delay $\tau$ is introduced between the end of the RF pulse and the onset of the DC pulse; negative delay $\tau<0$ corresponds to temporal overlap of the RF and DC pulses. We study both the $\tau\geq 0$ and $\tau<0$ regimes of RF-assisted switching. The phase of the RF pulse in our measurements is random (Supplementary Note 1 and Figure S1).
@@
Figure 5: RF-assisted increase in switching probability $\Delta\mathcal{P}$ as a function of RF frequency for $\tau=-2$ ns. Error bars are described in Supplementary Note 1.
@@
Figure 5 shows the dependence of $\Delta\mathcal{P}$ on $f_{\mathrm{RF}}$ for the $D=45$ nm device at $P_{\mathrm{RF}}=-21$ dBm (see Supplementary Note 1 and Table S1) and $\tau=-2$ ns. These data, taken at frequencies well below the FL FMR frequency (12 GHz, Supplementary Note 2 and Figure S3), reveal that RF-assisted switching becomes more efficient with decreasing RF frequency. A qualitatively similar dependence of $\Delta\mathcal{P}$ on $f_{\mathrm{RF}}$ is observed in the $\tau\geq 0$ regime.
Theoretical results
@@
In order to understand the mechanisms of the experimentally observed RF-assisted switching, we made room-temperature ($T=300$ K) numerical macrospin and micromagnetic simulations of the switching process. Simulations were made for all the MTJ diameters considered in the experiments (Supplementary Note 3). Here we show the results for the device with $D=45$ nm. Simulations of an ensemble of multiple replicas were performed to compute the switching probability. The magnetization dynamics of each replica were computed using both micromagnetic and macrospin solvers (see Supplementary Note 3). For the micromagnetic analysis, $10^{2}$ replicas were simulated, while for the macrospin simulations $10^{4}$ replicas were used. Both the micromagnetic and the macrospin simulations were performed using Magnetization Geometrical Integration Code (MaGICo, available at http://wpage.unina.it/mdaquino/index_file/MaGICo.html). The switching probability was then determined by counting the number of replicas that showed P to AP switching.
@@
For these simulations, we first select the values of $V_{\mathrm{DC}}$ and $t_{\mathrm{DC}}$ that produce $\mathcal{P}=0.5$ in the absence of an RF pulse for each device. We then add an RF pulse preceding the DC pulse and calculate the increase in the switching probability $\Delta\mathcal{P}$ above the $\mathcal{P}=0.5$ baseline. Figure 6 shows the simulated $\Delta\mathcal{P}$ due to the RF pulse assist versus normalized RF power $\tilde{P}_{\mathrm{RF}}$ and delay $\tau$ between the RF and DC write pulses. The RF power is normalized to the minimal theoretical power required to switch the magnetic state with a DC current in the absence of thermal fluctuations ($T=0$ K) [41]. This minimal power is analytically calculated in the macrospin approximation for the values of the FL dimensions, the uniaxial magnetocrystalline anisotropy constant $K_{\mathrm{AN}}$, saturation magnetization $M_{\mathrm{s}}$, and Gilbert damping constant $\alpha$ given in Table I of Methods.
Figure 6a shows macrospin $\Delta\mathcal{P}$ as a function of $\tilde{P}_{\mathrm{RF}}$ and $\tau$ when the phase constant $\phi$ of the RF voltage $\sqrt{2}V_{\mathrm{RF}}\cos(\omega_{\mathrm{RF}}t+\phi)$ is set to zero. Notice that $t=0$ corresponds to the time instant where the RF pulse is applied and $\omega_{\mathrm{RF}}=2\pi f_{\mathrm{RF}}$. These results are significantly different from the experiment in Fig. 3c and Fig. 4c. This is most clearly seen from the comparison of the inset in Fig. 6a showing $\Delta\mathcal{P}$ versus $\tau$ at $\tilde{P}_{\mathrm{RF}}=-5$ dB and the experimental data in Fig. 4c. Indeed, the measured $\Delta\mathcal{P}$ is a monotonically decreasing function of $\tau$, while the simulated $\Delta\mathcal{P}(\tau)$ has a peak at $\tau=-3$ ns. The micromagnetic results in Fig. 6b are similar to the macrospin with a slightly enhanced $\Delta\mathcal{P}$. The number of replicas in the macrospin simulations is large enough to yield a $\sim$1 % error in $\Delta\mathcal{P}$ while the micromagnetic results are precise within $\sim$10 % (Supplementary Note 3).
@@
The discrepancy with the experiment is expected because the simulations assume a fixed phase $\phi=0$, while the measurements average over random phases. To address this discrepancy, we performed macrospin simulations with $\phi$ uniformly distributed over $\pi/2,\pi/2]$ (Fig. 6c) and $\pi,\pi]$ (Fig. 6d) intervals, the latter directly mimicking the experimental conditions. The inset in Fig. 6d is qualitatively similar to the corresponding experimental results in Fig. 4c, although quantitative discrepancies remain. Similar results were obtained in micromagnetic simulations and for devices with other studied MTJ diameters (Supplementary Note 3 and Figure S4).
@@
(ii) prior to the RF+DC waveform, the distribution of replicas is assumed to be described by the Néel-Brown model [7] in the high-energy-barrier case.
@@
$\overline{V}_{\mathrm{ST}}=V_{\mathrm{DC}}+\sqrt{2}V_{\mathrm{RF}}\,\frac{|\tau|}{t_{\mathrm{DC}}}\mathrm{sinc}\left(\frac{\omega_{\mathrm{RF}}\tau}{2}\right)\cos\left(\omega_{\mathrm{RF}}\left(t_{\mathrm{RF}}+\frac{\tau}{2}\right)+\phi\right)\,$, (2)
which is $V_{\mathrm{DC}}$ plus the average RF voltage over the duration of the DC pulse, with $t_{\mathrm{RF}}$ the duration of the RF pulse. The theoretical justification of these five assumptions is discussed in Supplementary Note 3. Then, the magnetization switching dynamics is a damping-switching process in a uniaxial, bistable magnet between the stable states $m_{z0}$ and $m_{z1}$, where $m_{z0}$ ($m_{z1}$) belongs to the region of stability of $m_{z}=1$ ($m_{z}=-1$). The magnetization trajectory from $m_{z0}$ to $m_{z1}$ under the action of $\overline{V}_{\mathrm{ST}}$ can be expressed in closed form [15], which can be used to map the ensemble of initial conditions to the ensemble of final states. Therefore, the stochastic nature of the magnetization process is included only in the statistical distribution of the initial conditions. The FL switches if $m_{z}$ changes sign compared to $m_{z0}$ after the write pulse. Starting with the initial distribution with $m_{z0}>0$, the switching probability can be computed by integrating the PDF over the range $m_{z}\in[-1,0]$ that defines the well of the switched magnetic state $m_{z}=-1$. This PDF solution is described in Supplementary Note 3. The switching probability $\mathcal{P}$ for $t^{\prime}=t-(t_{\mathrm{RF}}+\tau)>0$ is given by:
[…]
The quantity $\overline{H}$ is the effective DC voltage pulse amplitude normalized to $V_{\mathrm{c0}}$, which is the critical DC voltage for switching at $T=0$. The quantity $\mu_{z}^{-1}(0,t^{\prime},\overline{H})$ is the initial condition $m_{z0}$ that by the deterministic dynamics under the action of the normalized voltage $\overline{H}$ for a time period of duration $t^{\prime}$ leads to the final state $m_{z1}=0$. The quantity $k_{\mathrm{eff}}/\tilde{T}$ is the ratio between the normalized effective anisotropy and the normalized temperature. They are given by: $k_{\mathrm{eff}}=2K_{\mathrm{AN}}/(\mu_{0}M_{s}^{2})-D_{\mathrm{z}}+D_{\perp}$, and $\tilde{T}=k_{\mathrm{B}}T/(\mu_{0}M_{s}^{2}V)$. Here, $K_{\mathrm{AN}}$ is the uniaxial magnetocrystalline anisotropy, $D_{\mathrm{z}}$ and $D_{\mathrm{\perp}}$ are the demagnetizing factors along the symmetry axis and the perpendicular direction $(2D_{\perp}+D_{z}=1)$, $k_{\mathrm{B}}$ is the Boltzmann constant, $\mu_{0}$ is the vacuum magnetic permeability, $M_{s}$ is the saturation magnetization, and $V$ is the volume of the free layer. All the quantities are expressed in SI units. Taking into account that PDF depends on $\phi$ via $\overline{H}$, the switching probability for $\phi$ uniformly distributed over the experimentally relevant interval $\pi,\pi]$ is:
@@
$\mathcal{P}(t^{\prime})=\frac{\int_{0}^{2\pi}\mathrm{erfi}\left(\sqrt{\frac{k_{\mathrm{eff}}}{2\tilde{T}}}\mu_{z}^{-1}(0,t^{\prime},\overline{H})\right)d\phi}{2\pi\,\mathrm{erfi}\left(\sqrt{\frac{k_{\mathrm{eff}}}{2\tilde{T}}}\right)}\,$. (4)
@@
In the case of the fully random phase $\phi$ distributed in $\phi\in[-\pi,\pi]$, analytical calculations in Fig. 7c do not predict an increase of $\mathcal{P}$, while numerical simulations in Fig. 6d show $\Delta\mathcal{P}>0$ for large values of $P_{\mathrm{RF}}$ and $\tau<0$. This is because $\mathcal{P}(\phi)$ in Eq. 3 is a periodic function of $\phi$, and the enhancement of $\mathcal{P}$ achieved for a given $\phi^{*}$ is approximately canceled by the reduction of $\mathcal{P}$ for the phase $\phi^{*}+\pi$. Although this cancellation is not exact, $\Delta\mathcal{P}$ resulting from the averaging in $\phi$ is small $\sim 0.01$.
In numerical simulations in Fig. 6d with sufficiently high $\tilde{P}_{\mathrm{RF}}$ and $\tau<0$, $\Delta\mathcal{P}$ is not negligibly small because a larger number of switching events with respect to those predicted by the analytical model in Fig. 7c occur in the time interval where the RF pulse adds constructively to the DC pulse voltage. This fact is not surprising since the assumption (v) of the analytical model does not hold anymore (Supplementary Note 3). As a result, numerical simulations produce higher $\Delta\mathcal{P}$ than the analytic model.